OpenStudy (anonymous):

-2(3x4)= -2x+2 please work out the prblem

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

idk

OpenStudy (skullpatrol):

Simplify the left side of the equation: -2(3x4)

OpenStudy (anonymous):

ok keeo going

OpenStudy (skullpatrol):

-2(3x4) = ?

OpenStudy (anonymous):

can you just work the problem out for me please

OpenStudy (skullpatrol):

Work "with" me please...

OpenStudy (anonymous):

im doing a test now thats why im asking

OpenStudy (skullpatrol):

-2(3x4) = -24

OpenStudy (anonymous):

thank you but they need the steps too

OpenStudy (skullpatrol):

-2(3x4)= -2(12) = -24 right?

OpenStudy (skullpatrol):

-2(3x4)= -2x+2 -2(12)= -2x+2 -24= -2x+2

OpenStudy (anonymous):

the answers the test had is a.2over -3 b.3over 2 c.2 d.24

OpenStudy (skullpatrol):

Do you want to learn how to do these type pf questions?

OpenStudy (anonymous):

after the test

OpenStudy (anonymous):

-2a+5a-4=11